# Spinal open books and symplectic fillings with exotic fibers

### v1

@@ 2.2. Monodromy factorizations
$\partial_{B}P:=\{c\in\partial P\mid c\text{ meets a spine component }\Sigma_{i}\times S^{1}\text{ such that }n_{i}\neq 0\}.$
@@ 2.4. Nearly Lefschetz fibrations
For a symplectic filling of an honest planar open book decomposition with suitable conditions, a pseudoholomorphic foliation gives a Lefschetz fibration [Wen10b]. However, symplectic fillings of a general spinal open book admit a generalized version of a Lefschetz fibration, which we call a nearly Lefschetz fibration. In particular, it gives rise to a uniform spinal open book on its boundary, which is non-Lefschetz-amenable in general.
@@
    For each $c\in B^{\operatorname{\text{exot}}}$, there are local holomorphic coordinates $\psi\colon U_{c}\to\mathbb{C}$, where $U_{c}\subset B^{\circ}$, sending $c$ to 00 and $\phi\colon V_{c}\to(\mathbb{C}-D)\times\mathbb{C}$, where $V_{c}\subset N(\partial_{h}(E))\cap\Pi^{-1}(U_{c})$ and $D$ is a small neighborhood of $0\in\mathbb{C}$, such that the following diagram commutes
@@
    $\Pi|_{\partial_{h}E\setminus\Pi^{-1}(B_{\operatorname{\text{exot}}})}:\partial_{h}E\setminus\Pi^{-1}(B_{\operatorname{\text{exot}}})\to B\setminus B_{\operatorname{\text{exot}}}$ is a smooth fiber bundle.
@@ 2.5. Positive multisections and spinal mapping classes
Let $\Pi\colon E\to B$ be a bordered Lefschetz fibration. A properly embedded compact oriented surface $\Sigma\subset E$ is called a positive multisection or a positive $n$–section of $E$ if it satisfies the following conditions:
@@
    The restriction $\Pi|_{\Sigma}$ is an $n$–fold simple branched covering map.
@@ 2.6. The relative ordinary open book of an annulus vertebra
[…]
@@ 2.7. Spine removal cobordisms
[LHMW18, Theorem D] Consider the contact manifold $(\widetilde{M},\widetilde{\xi})$, obtained from $(M,\xi)$ by deleting the spine component $\Sigma_{\text{rem}}\times S^{1}$ and capping off the corresponding boundary components of the pages by disks. Then there exists a symplectic cobordism with strongly concave boundary $(M,\xi)$ and weakly convex boundary $(\widetilde{M},\widetilde{\xi})$, defined by attaching $\Sigma_{\text{rem}}\times\mathbb{D}^{2}$ with a product symplectic structure along $\Sigma_{\text{rem}}\times S^{1}$.
@@ 3.1. Stable Hamiltonian structures and symplectic completions
A stable Hamiltonian structure on a $(2n-1)$-manifold is a pair $\mathcal{H}=(\Omega,\Lambda)$ where $\Omega$ is a $2$-form and $\Lambda$ is a $1$-form satisfying $\Lambda\wedge\Omega^{n-1}>0$, $d\Omega=0$ and $\ker\Omega\subset d\Lambda$.
Similar to contact structures, we may consider the hyperplane field $\Xi:=\ker\Lambda$ and associate to $\mathcal{H}$ a unique vector field $R_{\mathcal{H}}$, also called the Reeb vector field, given by $\Omega(R_{\mathcal{H}},\cdot)\equiv 0$ and $\Lambda(R_{\mathcal{H}})\equiv 1$. We may also define the symplectization of $(M,\mathcal{H})$ to be ×$\mathbb{R}\times M$ with the symplectic form
@@
where $\phi$ is any smooth function $\phi:\mathbb{R}\to(-\delta,\delta)$ with $\phi^{\prime}>0$ and $\phi(r)=r$ for $r$ near 00. The symplectic completion of a compact symplectic manifold $(W,\omega)$ with stable Hamiltonian boundary $\partial W=-M_{-}\sqcup M_{+}$ is the pair $(\widehat{W},\omega_{\phi})$ where
@@
As in the contact case, we consider compatible almost complex structure $J$ on $\mathbb{R}_{r}\times M$ with respect to the stable Hamiltonian structure $\mathcal{H}:=(\Omega,\alpha)$. This means that $J\partial r=R_{\mathcal{H}}$, $J(\Xi)=\Xi$ and $\Omega(\cdot,J\cdot)$ defines a bundle metric on $\Xi$. On the completion $(\widehat{W},\omega_{\phi})$, we consider the compatible almost complex structure $J$ to be such that $J|_{W}$ is compatible with $\omega$ and compatible with $\mathcal{H}_{-}$ and $\mathcal{H}_{+}$ on the symplectizations $(-\infty,0]\times M_{-}$ and $[0,\infty)\times M_{+}$. Note that any such compatible $J$ is $\omega_{\phi}$-tame on $\widehat{W}$ for every $\phi\in\mathcal{T}$. Therefore, the energy of a $J$-holomorphic curve $u:(S,j)\to(\mathbb{R}\times M,J)$ defined as
@@ 3.2. Moduli spaces of pseudoholomorphic foliations in a double completion
By construction, $\widehat{E}$ contains an unbounded closed subset $\widehat{N}_{+}(\partial E)\subset\widehat{E}$ which is identified with $[0,\infty)\times M^{+}$. Now, any $J_{+}$-compatible with $\mathcal{H}_{+}$ determines an $\omega_{E}$-compatible almost complex structure on $[0,\infty)\times M^{+}\cong\widehat{N}_{+}(\partial E)$. In [LHMW20, §3], a particular $J_{+}$ is carefully chosen so that there exists a $\mathbb{R}$-invariant foliation $\mathcal{F}_{+}$ on $[0,\infty)\times M^{+}$ that is $J_{+}$-holomorphic. The construction of the foliation $\mathcal{F}_{+}$ is in §3.8 in [LHMW20]. Again, we do not need the specific form of the foliation $\mathcal{F}_{+}$, but note that it includes “holomorphic pages” as shown in Proposition 3.12 in [LHMW20].
Now we discuss a specific model of the completion $(\widehat{W},\widehat{\omega})$ of a symplectic filling $(W,\omega)$ which is a variation of the construction in §3.1, but more suited for the discussion of pseudoholomorphic curves. Let $(M^{-},\xi_{-})$ be a weakly contact hypersurface obtained from $M^{0}$ by translation in the $s-$ and $t-$coordinates by a small negative value. In particular, $(M^{-},\xi_{-})$ is also contactomorphic to $(M,\xi)$. The completion is now given by
@@
where $\widehat{N}_{-}(\partial E)$ is the unbounded closed subset in $\widehat{E}$ with boundary $-M^{-}$ and the symplectic gluing is justified in [LHMW20, §6.1]. Note that $\widehat{N}_{-}(\partial E)\supset\widehat{N}_{+}(\partial E)$ and the almost complex structure $J_{+}$ constructed in the previous paragraph can be extended to an appropriate $\widehat{\omega}$-tame almost complex structure $\widehat{J}$ on $\widehat{W}$. Therefore, $(\widehat{W},\widehat{J})$ contains a cylindrical end $([0,\infty)\times M^{+},J_{+})$ compatible with $\mathcal{H}_{+}$.
Finally, we are ready to review the moduli spaces of interest defined in §6.1 of [LHMW20]. Let $\mathcal{M}(\widehat{J})$ be the moduli space of finite-energy holomorphic curves in $(\widehat{W},\widehat{J})$ with arithmetic genus zero and ¯$\overline{\mathcal{M}}(\widehat{J})$ be its compactification. Furthermore, define
^$\widehat{\mathcal{M}}(\widehat{J}):=\overline{\mathcal{M}}(\widehat{J})/\sim,$
where two curves are equivalent if and only if their asymptotic ends are the same up to permutation and their bottom-most nonempty levels are the same. It is easy to see that pseudoholomorphic leaves in the foliation $\mathcal{F}_{+}$ are elements in ^$\widehat{\mathcal{M}}(\widehat{J})$ as well.
In fact, identify $[0,\infty)\times M^{+}$ and ×$\mathbb{R}\times M^{+}$ in the usual way, and let ¯$\overline{\mathcal{M}}^{\mathcal{F}_{+}}_{2}(J_{+})$ denote the compactified moduli space of unparameterized index $2$ $J_{+}$-holomorphic curves in ×$\mathbb{R}\times M^{+}$ which belong to $\mathcal{F}_{+}$ modulo $\mathbb{R}$-translation, and
^$\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+}):=\overline{\mathcal{M}}^{\mathcal{F}_{+}}_{2}(J_{+})/\sim$
@@
^$\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+})\subset\widehat{\mathcal{M}}(\widehat{J}).$
@@
^$\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})\subset\widehat{\mathcal{M}}(\widehat{J})$
to be the smallest open and closed subset that contains ^$\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+})$. The main technical part of [LHMW20] is to give a partition of the interior of ^$\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})$ in terms of regular, singular and exotic buildings. Define ^$\widehat{\mathcal{M}}_{\operatorname{\text{reg}}}^{\mathcal{F}}(\widehat{J})$, ^$\widehat{\mathcal{M}}_{\operatorname{\text{sing}}}^{\mathcal{F}}(\widehat{J})$ and ^$\widehat{\mathcal{M}}_{\operatorname{\text{exot}}}^{\mathcal{F}}(\widehat{J})$ to consist of building whose main levels are curves in $\mathcal{M}_{\operatorname{\text{reg}}}(\widehat{J})$, $\mathcal{M}_{\operatorname{\text{sing}}}(\widehat{J})$ and $\mathcal{M}_{\operatorname{\text{exot}}}(\widehat{J})$, respectively, as in Theorem 1.1.
@@
^$\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})=\widehat{\mathcal{M}}_{\operatorname{\text{reg}}}^{\mathcal{F}}(\widehat{J})\cup\widehat{\mathcal{M}}_{\operatorname{\text{sing}}}^{\mathcal{F}}(\widehat{J})\cup\widehat{\mathcal{M}}_{\operatorname{\text{exot}}}^{\mathcal{F}}(\widehat{J})\cup\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+}),$
where ^$\widehat{\mathcal{M}}_{\operatorname{\text{reg}}}^{\mathcal{F}}(\widehat{J})$ is an open set and ^$\widehat{\mathcal{M}}_{\operatorname{\text{sing}}}^{\mathcal{F}}(\widehat{J})$ and ^$\widehat{\mathcal{M}}_{\operatorname{\text{exot}}}^{\mathcal{F}}(\widehat{J})$ are each a finite set. In addition, ^$\partial\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})=\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+})$ and every point in $\widehat{W}$ is the image of the main level of a unique curve in the interior of ^$\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})$ via the continuous surjection
@@ 3.3. Asymptotic properties of pseudoholomorphic foliations
In this subsection, we review the asymptotic properties of punctured pseudoholomorphic curves and behaviors of the moduli spaces described in §3.2. More details can be found in [HWZ96, HT07, Sie08]. Given a Reeb orbit $\gamma\subset Y$ and a pseudoholomorphic curve in ×$\mathbb{R}\times Y$ asymptotic to $\gamma$, one can naturally associate an asymptotic neighborhood ×$\mathbb{R}\times S^{1}\times\mathbb{R}_{\xi}^{2}$ to this asymptotic end, where the choice of $\mathbb{R}_{\xi}^{2}$ depends on a trivialization of the contact structure $\xi$.
@@
where #$\#\Gamma_{0}(u)$ denotes the number of asymptotic ends of $u$ with even Conley-Zehnder indices.
@@
This is an adaptation of the proof of Lemma A.2 in [Wen10b]. Since we are at an elliptic orbit, we may deform our almost complex structure via a path of admissible almost complex structure to $J^{\prime}$ such that $S(t)=\theta$, after further deforming the trivialization. See for example §2.6 of [HT07]. Notice that both choices of deformations are contractible.
Given $u\in\mathcal{M}$, recall that with respect to the splitting $u^{*}TW=T_{u}\oplus N_{u}$, we have a normal Cauchy–Riemann operator $D_{u}^{N}$ that is asymptotic to $L_{\gamma}$ and that $T_{u}\mathcal{M}=\ker D_{u}^{N}$. See [Wen10a, §3.4]. For a curve asymptotic to $\gamma$, one can choose coordinates $(s,t)\in[0,\infty)\times S^{1}$ near the puncture asymptotic to $\gamma$ so that for large $s,t$, we have the linearized asymptotic formula by by [HWZ96] (see also [Sie08]). Specifically, for any section $v\in\ker D_{u}^{N}$, we have that
@@
where $\lambda$ is an eigenvalue of $L_{\gamma}$, $f(t)$ is an eigenfunction of $L_{\gamma}$ with eigenvalue $\lambda$ and $r(s,t)\in\xi|_{\gamma}$ is smooth and converges uniformly to 00 in $t$ as $s\to\infty$.
@@ 4.1. A local model of exotic fibers
First, we identify the asymptotic neighborhood of an exotic fiber near the doubly-covered Reeb orbit with underlying simple orbit $\gamma_{e}$ with (\$\0(\mathbb{C}\backslash\{0\})\times\mathbb{C}$.
@@
The asymptotic neighborhood of a nondegenerate Reeb orbit $\gamma_{e}\subset(M,\xi)$, where $\xi$ is the stable Hamiltonian structure on $M$, admits a trivialization $\tau_{0}$ identified with ×$\mathbb{R}\times S^{1}\times\mathbb{R}^{2}$, with a local almost complex structure and a stable Hamiltonian structure inherited from the identification with (\$\0(\mathbb{C}\backslash\{0\})\times\mathbb{C}$ via the map
(7) $\begin{split}\phi:(\mathbb{C}\backslash\{0\})\times\mathbb{C}&\to\mathbb{R}\times S^{1}\times\mathbb{R}^{2}\\ \phi(r_{1}e^{i\theta_{1}},r_{2}e^{i\theta_{2}})&=(1/r_{1},-\theta_{1},r_{1}r_{2}e^{i(\theta_{1}+\theta_{2})})\end{split}$
@@
From [BEH+03] Example 3.1(4), we know that (\$(\mathbb{C}\backslash\{0\})\times\mathbb{C},i\oplus i)$ has cylindrical ends. The stable Hamiltonian structure associated to the level sets
$\0\{r_{1}e^{i\theta_{1}}\times\mathbb{C}|\theta_{1}\in S^{1}\}\subset(\mathbb{C}\backslash\{0\})\times\mathbb{C}$
@@
$\0u_{c}:\mathbb{C}\backslash\{\pm 1\}\to(\mathbb{C}\backslash\{0\})\times\mathbb{C}$
@@
(10) $\Pi_{0}:(\mathbb{C}\backslash\{0\})\times\mathbb{C}\to\mathbb{C}\text{ sending}(z_{1},z_{2})\mapsto z_{2}^{2}-z_{1},$

### v2

@@ 2.2. Monodromy factorizations
[…]
@@ 2.4. Nearly Lefschetz fibrations
For a symplectic filling of an honest planar open book decomposition with suitable conditions, a pseudoholomorphic foliation gives a Lefschetz fibration [Wen10c]. However, symplectic fillings of a general spinal open book admit a generalized version of a Lefschetz fibration, which we call a nearly Lefschetz fibration. In particular, it gives rise to a uniform spinal open book on its boundary, which is non-Lefschetz-amenable in general.
@@
    For each $c\in B^{\operatorname{\text{exot}}}$, there are local holomorphic coordinates $\psi\colon U_{c}\to\mathbb{C}$, where $U_{c}\subset B^{\circ}$, sending $c$ to 0 and $\phi\colon V_{c}\to(\mathbb{C}-D)\times\mathbb{C}$, where $V_{c}\subset N(\partial_{h}(E))\cap\Pi^{-1}(U_{c})$ and $D$ is a small neighborhood of $0\in\mathbb{C}$, such that the following diagram commutes
@@
    The restricted projection $\Pi|_{\partial_{h}E\setminus(\cup_{c\in B_{\operatorname{\text{exot}}}}V_{c})}:\partial_{h}E\setminus(\cup_{c\in B_{\operatorname{\text{exot}}}}V_{c})\to B\setminus\cup_{c\in B_{\operatorname{\text{exot}}}}\psi(V_{c})$ is a smooth fiber bundle, that extends to a smooth fiber bundle $\Pi_{h}:\partial_{h}E\to B$.
@@ 2.5. Positive multisections and spinal mapping classes
Let $\Pi\colon E\to B$ be a bordered Lefschetz fibration. A properly embedded compact oriented surface $\Sigma\subset E$ is called a positive multisection or a positive nn–section of $E$ if it satisfies the following conditions:
@@
    The restriction $\Pi|_{\Sigma}$ is an nn–fold simple branched covering map.
@@
Conversely, from a positive allowable monodromy factorization $t_{c_{1}}t_{c_{2}}\dots t_{c_{k}}\tau_{\gamma_{1}}\tau_{\gamma_{2}}\dots\tau_{\gamma_{m}}\in\operatorname{SMod}(F^{n})$, we can construct a positive allowable bordered Lefschetz fibration with a positive $n$-section.
@@ 2.6. The relative ordinary open book of an annulus vertebra
Van-Horn-Morris showed that $\Phi((aba)^{-4})=\text{id}$ and by inserting the word $(aba)^{-4}$, we can add $2\pi$-twisting to a torus bundle without changing the monodromy. Thus by inserting the word $(aba)^{-2}$, we can add $\pi$-twisting to a torus bundle, but change the monodromy by $\text{id}$. Without identifying the two ends of $T^{2}\times I$, the relative open book corresponding to the word $(aba)^{-2}$ gives the $\pi$-twisting contact $T^{2}\times I$, which is contactomorhpic to an annulus spine component. ∎
@@ 2.7. Spine removal cobordisms
[…]
@@ 3.1. Stable Hamiltonian structures and symplectic completions
A stable Hamiltonian structure on a (2n1)(2n-1)-manifold is a pair $\mathcal{H}=(\Omega,\Lambda)$ where $\Omega$ is a $2$-form and $\Lambda$ is a $1$-form satisfying $\Lambda\wedge\Omega^{n-1}>0$, $d\Omega=0$ and $\ker\Omega\subset d\Lambda$.
Similar to contact structures, we may consider the hyperplane field $\Xi:=\ker\Lambda$ and associate to $\mathcal{H}$ a unique vector field $R_{\mathcal{H}}$, also called the Reeb vector field, given by $\Omega(R_{\mathcal{H}},\cdot)\equiv 0$ and $\Lambda(R_{\mathcal{H}})\equiv 1$. We may also define the symplectization of $(M,\mathcal{H})$ to be $\mathbb{R}\times M$ with the symplectic form
@@
where $\phi$ is any smooth function $\phi:\mathbb{R}\to(-\delta,\delta)$ with $\phi^{\prime}>0$ and $\phi(r)=r$ for $r$ near 0. The symplectic completion of a compact symplectic manifold $(W,\omega)$ with stable Hamiltonian boundary $\partial W=-M_{-}\sqcup M_{+}$ is the pair $(\widehat{W},\omega_{\phi})$ where
@@
As in the contact case, we consider compatible almost complex structure $J$ on $\mathbb{R}_{r}\times M$ with respect to the stable Hamiltonian structure $\mathcal{H}:=(\Omega,\alpha)$. This means that $J\partial r=R_{\mathcal{H}}$, $J(\Xi)=\Xi$ and $\Omega(\cdot,J\cdot)$ defines a bundle metric on $\Xi$. On the completion $(\widehat{W},\omega_{\phi})$, we consider the compatible almost complex structure $J$ to be such that $J|_{W}$ is compatible with $\omega$ and compatible with $\mathcal{H}_{-}$ and $\mathcal{H}_{+}$ on the symplectizations $(-\infty,0]\times M_{-}$ and $0,\infty)\times M_{+}$. Note that any such compatible $J$ is $\omega_{\phi}$-tame on $\widehat{W}$ for every $\phi\in\mathcal{T}$. Therefore, the energy of a $J$-holomorphic curve $u:(S,j)\to(\mathbb{R}\times M,J)$ defined as
@@ 3.2. Moduli spaces of pseudoholomorphic foliations in a double completion
By construction, $\widehat{E}$ contains an unbounded closed subset $\widehat{N}_{+}(\partial E)\subset\widehat{E}$ which is identified with $0,\infty)\times M^{+}$. Now, any $J_{+}$-compatible with $\mathcal{H}_{+}$ determines an $\omega_{E}$-compatible almost complex structure on $0,\infty)\times M^{+}\cong\widehat{N}_{+}(\partial E)$. In [LHMW20, §3], a particular $J_{+}$ is carefully chosen so that there exists a $\mathbb{R}$-invariant foliation $\mathcal{F}_{+}$ on $0,\infty)\times M^{+}$ that is $J_{+}$-holomorphic. The construction of the foliation $\mathcal{F}_{+}$ is in §3.8 in [LHMW20]. Again, we do not need the specific form of the foliation $\mathcal{F}_{+}$, but note that it includes “holomorphic pages” as shown in Proposition 3.12 in [LHMW20].
[…]
@@
[…]
In fact, identify $0,\infty)\times M^{+}$ and $\mathbb{R}\times M^{+}$ in the usual way, and let $\overline{\mathcal{M}}^{\mathcal{F}_{+}}_{2}(J_{+})$ denote the compactified moduli space of unparameterized index 22 $J_{+}$-holomorphic curves in $\mathbb{R}\times M^{+}$ which belong to $\mathcal{F}_{+}$ modulo $\mathbb{R}$-translation, and
$\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+}):=\overline{\mathcal{M}}^{\mathcal{F}_{+}}_{2}(J_{+})/\sim$
@@
[…]
@@
[…]
@@
$\widehat{\mathcal{M}}^{\mathcal{F}}(\widehat{J})=\widehat{\mathcal{M}}_{\operatorname{\text{reg}}}^{\mathcal{F}}(\widehat{J})\cup\widehat{\mathcal{M}}_{\operatorname{\text{sing}}}^{\mathcal{F}}(\widehat{J})\cup\widehat{\mathcal{M}}_{\operatorname{\text{exot}}}^{\mathcal{F}}(\widehat{J})\cup\widehat{\mathcal{M}}^{\mathcal{F}_{+}}(J_{+}),$
[…]
@@ 3.3. Asymptotic properties of pseudoholomorphic foliations
In this subsection, we review the asymptotic properties of punctured pseudoholomorphic curves and behaviors of the moduli spaces described in §3.2. More details can be found in [HWZ96, HT07, Sie08]. Given a Reeb orbit $\gamma\subset Y$ and a pseudoholomorphic curve in $\mathbb{R}\times Y$ asymptotic to $\gamma$, one can naturally associate an asymptotic neighborhood $\mathbb{R}\times S^{1}\times\mathbb{R}_{\xi}^{2}$ to this asymptotic end, where the choice of $\mathbb{R}_{\xi}^{2}$ depends on a trivialization of the contact structure $\xi$.
@@
where $\Gamma_{0}(u)$ denotes the number of asymptotic ends of $u$ with even Conley-Zehnder indices.
First, we recall “automatic transversality” in dimension four, which guarantees transversality for certain curves without genericity assumption on the almost complex structures. The general version is in [Wen10a]. However, we only need the following special case.
Theorem 3.4.
[Wen10c, Proposition A.1] If $u$ is an immersed finite energy pseudoholomorphic curve with $\operatorname{ind}(u)>c_{N}$, then $u$ is unobstructed.
@@
This is an adaptation of the proof of Lemma A.2 in [Wen10c]. Since we are at an elliptic orbit, we may deform our almost complex structure via a path of admissible almost complex structure to $J^{\prime}$ such that $S(t)=\theta$, after further deforming the trivialization. See for example §2.6 of [HT07]. Notice that both choices of deformations are contractible and that local diffeomorphisms are preserved under these deformations.
Given $u\in\mathcal{M}$, recall that with respect to the splitting $u^{*}TW=T_{u}\oplus N_{u}$, we have a normal Cauchy–Riemann operator $D_{u}^{N}$ that is asymptotic to $L_{\gamma}$ and that $T_{u}\mathcal{M}=\ker D_{u}^{N}$. See [Wen10a, §3.4]. By Theorem 3.4, we also know that $\dim\ker D_{u}^{N}=2$. For a curve asymptotic to $\gamma$, one can choose coordinates $(s,t)\in[0,\infty)\times S^{1}$ near the puncture asymptotic to $\gamma$ so that for large $s,t$, we have the linearized asymptotic formula by by [HWZ96] (see also [Sie08]). Specifically, for any section $v\in\ker D_{u}^{N}$, we have that
@@
where $\lambda$ is an eigenvalue of $L_{\gamma}$, $f(t)$ is an eigenfunction of $L_{\gamma}$ with eigenvalue $\lambda$ and $r(s,t)\in\xi|_{\gamma}$ is smooth and converges uniformly to 0 in $t$ as $s\to\infty$.
@@ 4.1. A local model of exotic fibers
[…]
@@
[…]
@@
[…]
@@
$u_{c}:\mathbb{C}\backslash\{\pm 1\}\to(\mathbb{C}\backslash\{0\})\times\mathbb{C}$
@@
[…]